OpenStudy (anonymous):

How do you know if you have to factor a quadratic expression to solve a rational equation?

OpenStudy (anonymous):

You don't ;D

hartnn (hartnn):

For, \(ax^2+bx+c\) If discriminant \(D = b^2-4ac\) is a PERFECT SQUARE only then, you can factor your quadratic expression into integral factors.

OpenStudy (anonymous):

Can you show me with this problem? 1/t-2=t/8

hartnn (hartnn):

is "t-2" in the denominator ?? then multiply throughout by t-2 if only "t" is in the denominator, then multiply throughout by 't'

OpenStudy (anonymous):

The answers are -2,4 . But how do they get that?

hartnn (hartnn):

so its (t-2) 1/(t-2)=t/8

hartnn (hartnn):

multiply both sides by 8(t-2) 8 = t (t-2) got this ?

OpenStudy (anonymous):

Yes

hartnn (hartnn):

so t^2 -2t -8 = 0 do you know how to solve this quadratic ?

OpenStudy (anonymous):

Still confused, Can you write it out step by step?

hartnn (hartnn):

you never solved a quadratic before ?? can you tell me 2 numbers whose sum is -2 and product is -8 ?

OpenStudy (anonymous):

-4 and 2?

hartnn (hartnn):

correct! so we split -2t as \(\Large -4t+2t\)

hartnn (hartnn):

\(x^2-4t+2t -8 =0 \) factor out x from first 2 terms and 2 from the last 2 terms

hartnn (hartnn):

\(t(t-4) +2(t-4) =0\)

hartnn (hartnn):

\((t-4)(t+2) = 0\)

hartnn (hartnn):

t-4 = 0 or t+2 =0 t=4 or t = -2

OpenStudy (anonymous):

I got it!!! :D Thank you kind stranger!

hartnn (hartnn):

welcome ^_^ happy to help :)